HETAV PATEL

Grade 11,

How to substute values of 25a*85d when a=2 and d=5 and also50a-4m when a=70 and m=30

How to substute values of 25a*85d when a=2 and d=5 and also50a-4m when a=70 and m=30

Grade:7

2 Answers

Arun
4 years ago
Dear student
 
In 25 a *85d
 
Put a = 2 and d = 5
 
25 *2*85*5 = 50 * 425 = 21250
 
Hope it helps
 
PS- please ask only one question in one session
Sayantan Garai
4 years ago
As  a=2 and d=5,
25a*85d = 25 * 2 * 85 * 5 =50 * 425 = 21250.
 
Also, 50a-4m = 50*10 - (4*30)   [a=10 and m =30]
                             = 500 - 120 = 380.